before colliding, the momentum of block A is +25.0 kg*m/s, and block b is +35.0 kg*m/s. after, block a has a momentum +60.0 kg*m/s. what is the momentum of block b afterwards?
Answer #1

| Pa, = 25 kg-m, Pb, = 35 kg-m nivena) í Paz = 60 kg-mg 12 Apply momentum conservation, Pa, tPb, = Paz + Pb2 25+35 = 60 + Pb2